# White Dwarf Structure and Binary Inspiral Gravitational Waves from Quantum Hadrodynamics

### v1

@@ 3 The structure of white dwarfs and binary inspirals
where $\mathcal{N}$ is density of the atom-like unit number. It should be noted that the pure mesonic parts of energy density disappear outside the nuclei. Our numerical results of nucleon density distributions of different nuclei and the corresponding EOS of WD matter are shown in App. A.
3.1 Structure of white dwarf
After obtaining $\mathcal{E}-P$ relation, the M-R relations can be calculated by solving the Tolman-Oppenheimer-Volkoff (TOV) equations Tolman (1939); Oppenheimer & Volkoff (1939) and the TD of WDs can also be obtained following Ref. Flanagan & Hinderer (2008). It should be noted that in our present work, we will not concern about the effect of $\beta$-equilibrium Salpeter (1961); Rotondo et al. (2011b, a); Boshkayev et al. (2013). Our results are shown in Fig. 2, where it can be seen that our WD structures saturate the regular Chandrasekhar limit (free electron gas, a pivotal concept in astrophysics). A comparison of the pressure at different mass density obtained from our approach, Chandrasekhar limit, Salpeter approach, and the relativistic FMT from Ref. Rotondo et al. (2011a) are shown in Table 3. It can be seen that, in our approach, the structure of WDs is consistent with the Salpeter approach and the relativistic FMT framework, which slightly deviates from the Chandrasekhar limit. Our results align well within the region of the observational data, thereby validating the effectiveness of our theoretical framework in describing the properties of WDs.
{widetext}
@@ 3.1 Structure of white dwarf
Table 3: The pressure at different mass densities $\rho_{\rm mass}$ for regular Chandrasekhar limit ($P_{\rm Ch}$), Salpeter approach ($P_{\rm S}$), the relativistic FMT ($P_{\rm FMT}^{\rm rel}$), our parametrization with free electron gas limit ($P_{\rm RMF}^{\rm Ch}$), and our parametrization ($P_{\rm RMF}$). $\rho_{\rm mass}$ is unit of $\rm g/cm^{3}$, and pressure is in the unit of $\rm dyne/cm^{2}$. We take 12C as an example.
$\rho_{\rm mass}$ $P_{\rm Ch}$ $P_{\rm S}$ $P_{\rm FMT}^{\rm rel}$ $P_{\rm RMF}^{\rm Ch}$ $P_{\rm RMF}$
$10^{4}$   $1.45\times 10^{19}$   $1.29\times 10^{19}$  $1.29\times 10^{19}$  $1.45\times 10^{19}$  $1.29\times 10^{19}$
$10^{5}$   $6.50\times 10^{20}$   $6.14\times 10^{20}$  $6.13\times 10^{20}$  $6.48\times 10^{20}$  $6.13\times 10^{20}$
$10^{6}$   $2.63\times 10^{22}$   $2.55\times 10^{22}$  $2.54\times 10^{22}$  $2.62\times 10^{22}$  $2.54\times 10^{22}$
$10^{7}$   $8.46\times 10^{23}$   $8.29\times 10^{23}$  $8.27\times 10^{23}$  $8.44\times 10^{23}$  $8.28\times 10^{23}$
$10^{8}$   $2.15\times 10^{25}$   $2.11\times 10^{25}$  $2.11\times 10^{25}$  $2.14\times 10^{25}$  $2.11\times 10^{25}$
$10^{9}$   $4.86\times 10^{26}$   $4.78\times 10^{26}$  $4.77\times 10^{26}$  $4.84\times 10^{26}$  $4.77\times 10^{26}$
$10^{10}$   $1.06\times 10^{28}$   $1.04\times 10^{28}$  $1.04\times 10^{28}$  $1.05\times 10^{28}$  $1.04\times 10^{28}$
@@
Figure 2: The M-R relation (left panel) and TD (right panel) of WDs. The observational data marked by "$\times$" are taken from Ref. Tremblay et al. (2016). The Chandrasekhar limit of our parametrization (denoted as ’Ch’) is also plotted for comparison.
3.2 Gravitational waves of binary inspirals
@@
We finally calculate the GW signals from the binary WD inspirals with and without tidal deformability at the 2.5PN order. The expression of the GW waveforms up to 2.5PN order are listed in the App. B. The results of the GWs from binary WDs composed of different elements in frequency-domain, are shown in Fig. 3 and Fig. 4 for, respectively, the phase shift and amplitude deviation. We typically choose the equal mass system with component mass $1.1\ M_{\odot}$. The corresponding tidal deformability are Λ=7352, 6613, $\Lambda=7352,\ 6613,\ 6197$ for ${}^{4}\rm He$, ${}^{12}\rm C$, and ${}^{16}\rm O$, respectively, according to the results of Fig. 2.
@@
Figure 3: The phase shift of GWs in mHz region (left panel) and hundred Hz region (right panel).
@@
Figure 4: The amplitude deviation of GWs in mHz region (left panel) and hundred Hz region (right panel) introduced by TD.
@@
Our results illustrated Fig. 3 and Fig. 4 indicate that WDs composed of lighter elements show a bigger deviation from the point-particle approximation—the GW signals without TD—in terms of both phase and amplitude, in line with the tidal deformability shown in Fig. 2. Although the PN approximation may not be suitable for the merging stage of WD inspirals, particularly in the hundred Hz frequency range, the discernible differences between the point particle approximation and the tidal deformability corrections offer a preliminary understanding of matter effects on GWs. This is especially valuable given the complexities involved in numerical simulations for such systems. Comparative analyses in both the mHz and hundred Hz regions could serve as initial guides for future GW detections and as a testbed for further research on WDs. Moreover, understanding the point particle approximation’s limitations in GW physics is essential for interpreting GW signals accurately.
Our results of the dimensionless characteristic strain amplitude $h_{c}(f)$ Moore et al. (2015)
@@
in comparison with the detectabilities of LISA Belgacem et al. (2019), Taiji Luo et al. (2020, 2021), and TianQin Gong et al. (2021), at $r=$3Mpc and 300Mpc, are presented in Fig. 5. One can see that, as expected, the events of the inspirals at $\lesssim 300$Mpc are indeed detectable by these space-based facilities.
Figure 5: The predicted characteristic strain amplitude of GWs in comparision with the detectabilities of space-based facilities.
4 Summary and perspective
In this work, we successfully extended the RMF approach to study the WDs and proposed an atom-like constituent of WDs. The achieved properties of nuclei and WDs are closely aligned with observational data. The GWs stemming from the merger of binary WDs are also studied. A notable advantage of our approach is the ability to describe NSs, NS crusts, and WDs within a single, cohesive framework by employing distinct BCs.
Upon the parameters established from the properties of ordinary homogeneous NM at saturation density, we proceeded to determine the structure of nuclei. The results obtained are closely aligned with empirical data Fortune (2016); National Nuclear Data Center (2024); Pössel (2010), except for the anomalously deep bounded ${}^{4}\rm He$ and nuclei with large neutron-proton ($N-Z$) differences, which are attributed to the simplified interaction parametrizations. Subsequently, we computed the M-R relations and TD of WDs conceptualizing them as an atom-like unit composed objects and found that the yielded results are in concordance with observational data Tremblay et al. (2016). This congruence underscores the efficacy of our model in capturing the essential physics of WDs.
Finally, we generated GW signals for binary WD inspirals composed of various elements, ${}^{4}\rm He$, ${}^{12}\rm C$, and ${}^{16}\rm O$, considering both scenarios with and without TD across mHz and hundred Hz frequency ranges, up to the 2.5 PN order. The resulting signals indicate that WDs composed of lighter elements exhibit bigger deviations from the point particle approximation, consistent with the TD findings. These results provide valuable preliminary guidelines for interpreting future GW detection signals.
@@ 4 Summary and perspective
Firstly, as well known, the present used Walecka-type model is not enough for precisely describe nuclei and NM. So that it is interesting to fine the calculation using more realistic models anchored on the chiral symmetry of QCD Jenkins & Manohar (1991); Bernard (2008); Scherer (2010); Ma & Ma (2024); Li et al. (2017); Ma & Rho (2017), ab initio parameterized nuclear force Bender et al. (2003); Stoitsov et al. (2010); Hergert et al. (2016); Duguet et al. (2015) or Skyme force parameterization Agrawal et al. (2006); Lesinski et al. (2006); Zuo et al. (2018). We expect these studies can advance our grasp of the interplay between nuclear structure physics and QCD and access to nuclei structure more precisely.
Secondly, so far we suppose that the WDs are made up of the atom-like units with a fixed nuclei. To take the shell structure of WDs into account, the transition between elements should be properly considered in the future work.
Thirdly, although this preliminary work catches the dominant physics of WD, we only considered the WD made of a unique nucleus. It is accepted that the inverse beta-decay of nucleus plays a fundamental role in WD physics Rotondo et al. (2011b, a); Boshkayev et al. (2013). This deserves a serious consideration in the present framework.
@@ Appendix A The nuclei structures and EOS of white dwarfs
The GW in the frequency-domain ($f$) can be obtained via Husa et al. (2016); Isoyama et al. (2020),
$\tilde{h}(f)$ $=$ $\int h(t)e^{i2\pi ft}\mathrm{~{}d}t=A(f)e^{i\left(\psi_{\operatorname{SPA}}(f)-\pi/4\right)},$ (B1)
where $\psi_{\mathrm{SPA}}(f)=2\pi ft(f)-\psi(f)$. $\psi(f)$ can be decomposed into point and tidal parts as $\psi(f)=\psi_{\mathrm{pp}}(f)+\psi_{\text{tidal }}(f)$ and the same treatment can be applied to amplitude $A(f)=A_{\mathrm{pp}}(f)+A_{\text{tidal }}(f)$. The explicit expressions for the amplitude and phase are Isoyama et al. (2020); Buonanno et al. (2009); Kawaguchi et al. (2018)
$A_{\rm pp}(f)$ $\simeq$ $\frac{\mathcal{M}^{5/6}}{r}\sqrt{\frac{2}{3\pi^{1/3}}}f^{-7/6}\left(1+O\left(f^{2/3}\right)\right)\ ,$ (B2)
@@
$A_{\text{tidal }}(f)=\sqrt{\frac{5\pi\nu}{24}}\frac{M^{2}}{r}\tilde{\Lambda}v^{-7/2}\left(-\frac{27}{16}v^{10}-\frac{449}{64}v^{12}-4251v^{15.780}\right),$ (B3)
@@
$\psi_{\mathrm{pp}}(f)$ $=$ $2\pi ft_{c}-\phi_{c}-\frac{\pi}{4}+\frac{3}{128\nu v^{5}}\left[1+\frac{20}{9}\left(\frac{743}{336}+\frac{11}{4}\nu\right)v^{2}-16\pi v^{3}\right.$ (B4)
@@
$\psi_{\text{tidal }}(f)=\frac{3v^{5}}{128\nu}\left[-\frac{39}{2}\tilde{\Lambda}\left(1+12.55\tilde{\Lambda}^{2/3}v^{8.480}\right)\right]\times\left(1+\frac{3115}{1248}v^{2}-\pi v^{3}+\frac{28024205}{3302208}v^{4}-\frac{4283}{1092}\pi v^{5}\right)\ .$ (B5)

### v2

@@ III Structure of white dwarfs and binary inspirals
It should be noted that the pure mesonic contributions to the energy density vanish outside the nuclei, while the electron density disappears within the nucleus. Our numerical results of the nucleon density distributions of different nuclei and the corresponding EOS of WD matter are presented in App. A.
III.1 Structure of white dwarf
After obtaining $\mathcal{E}-P$ relation, the M-R relations can be calculated by solving the Tolman-Oppenheimer-Volkoff (TOV) equations [60, 44] and the TD of WDs can also be obtained by solving the Love number equation following Ref. [17, 46]. It should be noted that in our current work, we will not consider the effect of $\beta$-equilibrium [51, 48, 49, 11].
Our results are resented in Fig. 2, where it can be seen that our WD structures saturate the regular Chandrasekhar limit (free electron gas, a pivotal concept in astrophysics) [13]. A comparison of the pressure at different mass density obtained from our full calculation ($P_{\rm RMF}$) and free electron limit ($P_{\rm RMF}^{\rm Ch}$), Chandrasekhar limit ($P_{\rm Ch}$[13], Salpeter approach ($P_{\rm S}$[51], and the relativistic FMT from Ref. [49] ($P_{\rm FMT}^{\rm rel}$) is shown in Table 3. It can be seen that, in our approach, the EOS of WDs deviates slightly more from the Chandrasekhar limit compared to the Salpeter approach and the relativistic FMT framework, both of which provide corrections to the WD structures at fine structure level.
@@ III.1 Structure of white dwarf
Table 3: The pressure at different mass densities $\rho_{\rm mass}$ for regular Chandrasekhar limit ($P_{\rm Ch}$[13], Salpeter approach ($P_{\rm S}$[51], the relativistic FMT ($P_{\rm FMT}^{\rm rel}$[49], our parametrization with free electron gas limit ($P_{\rm RMF}^{\rm Ch}$), and our parametrization ($P_{\rm RMF}$). $\rho_{\rm mass}=E\cdot\rho_{\rm atoms}$ is unit of $\rm g/cm^{3}$, where $E$ is the total energy of every atomic unit and $\rho_{\rm atoms}$ is the number density of atom-like units, and pressure is in the unit of $\rm dyne/cm^{2}$. We take 12C as an example.
$\rho_{\rm mass}$ $P_{\rm Ch}$ [13] $P_{\rm S}$ [51] $P_{\rm FMT}^{\rm rel}$ [49] $P_{\rm RMF}^{\rm Ch}$ $P_{\rm RMF}$
$10^{4}$   $1.45\times 10^{19}$   $1.29\times 10^{19}$  $1.29\times 10^{19}$  $1.45\times 10^{19}$  $1.27\times 10^{19}$
$10^{5}$   $6.50\times 10^{20}$   $6.14\times 10^{20}$  $6.13\times 10^{20}$  $6.48\times 10^{20}$  $6.06\times 10^{20}$
$10^{6}$   $2.63\times 10^{22}$   $2.55\times 10^{22}$  $2.54\times 10^{22}$  $2.62\times 10^{22}$  $2.52\times 10^{22}$
$10^{7}$   $8.46\times 10^{23}$   $8.29\times 10^{23}$  $8.27\times 10^{23}$  $8.44\times 10^{23}$  $8.20\times 10^{23}$
$10^{8}$   $2.15\times 10^{25}$   $2.11\times 10^{25}$  $2.11\times 10^{25}$  $2.14\times 10^{25}$  $2.09\times 10^{25}$
$10^{9}$   $4.86\times 10^{26}$   $4.78\times 10^{26}$  $4.77\times 10^{26}$  $4.84\times 10^{26}$  $4.74\times 10^{26}$
$10^{10}$   $1.06\times 10^{28}$   $1.04\times 10^{28}$  $1.04\times 10^{28}$  $1.05\times 10^{28}$  $1.03\times 10^{28}$
@@
Figure 2: The M-R relation (upper panel) and TD (lower panel) of WDs. The observational data marked by $\times"$ are taken from Ref. [62]. The Chandrasekhar limit of our parametrization (denoted as ’Ch’) is also plotted for comparison.
We compare in Fig. 3 the M-R relations calculated from our model (RMF) and relativistic FMT [49], Salpeter approach [51] and Chandrasekhar limit [13].
Figure 3: Comparison of the M-R relations of $^{12}\rm C$ WD calculated from our model (RMF), relativistic FMT [49], Salpeter approach [51] and Chandrasekhar limit [13]. It should be noted that the M-R relation of relativistic FMT is calculated by using higher order TOV equations, while others are calculated by using the standard TOV equations.
Our results align well within the region of the observational data, thereby validating the effectiveness of our theoretical framework in describing the properties of WDs. Furthermore, a realistic description of the nuclei within WDs will yield corrections of a similar magnitude to those arising from considerations of electromagnetic interactions or relativistic fluid corrections in WD matter. This signifies the importance of including the nucleus structure in the WD study.
III.2 Gravitational waves of binary inspirals
@@
We finally calculate the GW signals from the binary WD inspirals. In Ref. [71], the GW signals emitted from WD coalescences were studied in the Newtonian limit. Here, we will investigate the signals with and without tidal deformability at the 2.5PN order. The expression of the GW waveforms up to the 2.5PN order are given in App. B. The results of the GWs from binary WDs composed of different elements in frequency-domain, are shown in Fig. 4 and Fig. 5 for, respectively, the phase shift and amplitude deviation. We typically choose the equal mass system with component mass $1.1\ M_{\odot}$. The corresponding tidal deformability are $\Lambda=7352,\ 6613,\ 6197$ for $^{4}\rm He$, $^{12}\rm C$, and $^{16}\rm O$, respectively, according to the results shown in Fig. 2.
@@
Figure 4: The phase shift of GWs in mHz region (upper panel) and hundred Hz region (lower panel).
@@
Figure 5: The amplitude deviation of GWs in mHz region (upper panel) and hundred Hz region (lower panel) introduced by TD.
@@
Our results, as illustrated in Fig. 4 and Fig. 5, indicate that WDs composed of lighter elements exhibit a bigger deviation from the point-particle approximation—the GW signals without TD—in terms of both phase and amplitude, in line with the tidal deformability shown in Fig. 2. Although the PN approximation may not be suitable for the merging stage of WD inspirals, particularly in the hundred Hz frequency range, the discernible differences between the point particle approximation and the tidal deformability corrections offer a preliminary understanding of matter effects on GWs. This insight is especially valuable given the complexities involved in numerical simulations for such systems. Comparative analyses in both the mHz and hundred Hz regions could serve as initial guides for future GW detections and as a testbed for further research on WDs. Additionally, understanding the limitation of point particle approximation in GW physics is crucial for interpreting GW signals.
In addition, it is interesting to study the dimensionless characteristic strain amplitude $h_{c}(f)$ [41]
@@
where $S_{h}(f)$ is the power spectral density (PSD). Our results in comparison with the detectabilities of LISA [8], Taiji [36, 37], and TianQin [22], at $r=3$ Mpc and 300 Mpc, are presented in Fig. 6. With the result of characteristic strain $h_{c}(f)$, we calculate the signal-to-noise ratio (SNR) $\varrho$ defined as [41, 69]
$\varrho=\sqrt{\int^{f_{\max}}_{f_{\min}}\frac{4|\tilde{h}(f)|^{2}}{S_{n}(f)}\mathrm{d}f}=\sqrt{\int^{\log f_{\max}}_{\log f_{\min}}\left(\frac{h_{c}(f)}{h_{n}(f)}\right)^{2}\mathrm{d}(\log f)},$ (11)
where $S_{n}$ is the one-sided noise PSD, and $h_{n}(f)=\sqrt{fS_{n}(f)}$. In the frequency range 0.0420-0.0428 Hz within the Taiji’s mission duration $\sim 5$ years, $\varrho\sim 10$ for $r=3$ Mpc and $\varrho\sim 0.1$ for $r=300$ Mpc. One can see that, as expected, the events of the inspirals at $\lesssim 3$ Mpc are indeed detectable by these space-based facilities.
Figure 6: The predicted characteristic strain amplitude of GWs in comparison with the detectabilities of space-based facilities.
IV Summary and perspective
In this work, we successfully extend the RMF approach to study the WDs by using a Walecka-type model that incorporates hadrons, electrons and photons in a unified framework. We propose that a WD is composed universally of a single element atom-like constituents—Wigner-Seitz cells. The properties of the nuclei and WDs obtained from our model are closely aligned with observational data. Additionally, we investigate the GWs stemming from the merger of binary WDs.
Although estimates of nucleus structure or corresponding physical process in WD studies can be simply obtained through experimental data, we believe that the approach established in this work is particularly valuable for studing WDs composed of elements, whose structures are not well measured. Additionally, our framework is usefully for examining the process of continuous neutronization (beta equilibrium) in WDs and NS crusts with the density changing. That’s because an advantage of our approach is its ability to describe NSs (homogeneous nuclear matter), NS crusts (cluster of nuclear matter), and WDs (nucleus surrounded by electron gas) within a single, cohesive framework by employing distinct BCs.
Based on the parameters established from the properties of ordinary homogeneous NM at saturation density, we estimated the structure of nuclei. The results obtained are closely aligned with experiment data [18, 42], except for the anomalously deep bounded $^{4}\rm He$ and nuclei with large neutron-proton (NZN-Z) differences. This discrepancies are attributed to the simplified interaction parametrizations and ignorance of the shell effect. Subsequently, we computed the M-R relations and TD of WDs, conceptualizing them as an atom-like unit composed objects. We found that the yielded results are in concordance with observational data [62]. Our approach incorporates the size effect of nucleus, Coulomb force and photon energy in a unified model. The obtained results show a more obvious fine structure deviation compared to other methods. This congruence underscores the efficacy of our model in contributing to the understanding of the WD physics at a fine structure level.
Finally, we generated GW signals for binary WD inspirals composed of various elements, including $^{4}\rm He$, $^{12}\rm C$, and $^{16}\rm O$, considering both scenarios with and without TD across frequency ranges from mHz to hundred Hz, up to the 2.5 PN order. The resulting signals indicate that WDs composed of lighter elements exhibit bigger deviations from the point-particle approximation, which is consistent with the finding related to TD. These results provide valuable preliminary guidelines for interpreting future GW detection signals.
@@ IV Summary and perspective
Firstly, as is well known, the present used Walecka-type model is insufficient for accurately describing nuclei and NM. Therefore, it would be helpful and interesting to refine the calculations using more realistic models based on the chiral symmetry of QCD [29, 10, 53, 38, 34, 39], ab initio parameterized nuclear force [9, 57, 24, 16] or Skyrme force parametrization [4, 33, 72]. We expect that these studies will enhance our understanding of the interplay between nuclear structure physics and QCD, allowing us to access nuclear structure with greater precision.
Secondly, we have so far assumed that WDs are composed of the atom-like units with a fixed nucleus. To take the shell structure of WDs into account, the transition between elements should be properly considered in the future work.
Thirdly, although this preliminary work captures the dominant physics of WD, we have only considered WDs composed of a unique nucleus. It is accepted that the inverse beta-decay of nucleus plays a fundamental role in WD physics [48, 49, 11]. This aspect deserves a serious consideration within the current framework.
Last but not least, since binary WDs may get crystallized after a long evolution, it is interesting to consider how the crystallization of WDs affects their tidal deformations. Such changes could also contribute to the GWs generated during the inspiral and merger of binary WDs or binary white dwarf-neutron star systems. These phenomena are detectable by space-based facilities as discussed in [59, 45]. We leave this topic to our future work.
@@ Appendix A The nuclei structures and EOS of white dwarfs
The slight difference between distributions of neutrons and protons in $^{12}{\rm C}$ is due to EM interactions between protons.
@@
The GW in the frequency-domain ($f$) can be obtained via [27, 28],
$\tilde{h}(f)$ $=$ $\int h(t)e^{i2\pi ft}\mathrm{~d}t=A(f)e^{i\left(\psi_{\operatorname{SPA}}(f)-\pi/4\right)},$ (12)
where $\psi_{\mathrm{SPA}}(f)=2\pi ft(f)-\psi(f)$. $\psi(f)$ can be decomposed into point and tidal parts as $\psi(f)=\psi_{\mathrm{pp}}(f)+\psi_{\text{tidal }}(f)$. The same treatment applies to amplitude $A(f)=A_{\mathrm{pp}}(f)+A_{\text{tidal }}(f)$. The explicit expressions for the amplitude and phase are [28, 12, 30]
$A_{\rm pp}(f)$ $\simeq$ $\frac{\mathcal{M}^{5/6}}{r}\sqrt{\frac{2}{3\pi^{1/3}}}f^{-7/6}\left(1+O\left(f^{2/3}\right)\right)\ ,$ (13)
@@
$A_{\text{tidal }}(f)=\sqrt{\frac{5\pi\nu}{24}}\frac{M^{2}}{r}\tilde{\Lambda}v^{-7/2}\left(-\frac{27}{16}v^{10}-\frac{449}{64}v^{12}-4251v^{15.780}\right),$ (14)
@@
$\psi_{\mathrm{pp}}(f)$ $=$ $2\pi ft_{c}-\phi_{c}-\frac{\pi}{4}+\frac{3}{128\nu v^{5}}\left[1+\frac{20}{9}\left(\frac{743}{336}+\frac{11}{4}\nu\right)v^{2}-16\pi v^{3}\right$. (15)
@@
$\psi_{\text{tidal }}(f)=\frac{3v^{5}}{128\nu}\left[-\frac{39}{2}\tilde{\Lambda}\left(1+12.55\tilde{\Lambda}^{2/3}v^{8.480}\right)\right]\times\left(1+\frac{3115}{1248}v^{2}-\pi v^{3}+\frac{28024205}{3302208}v^{4}-\frac{4283}{1092}\pi v^{5}\right)\ .$ (16)